harp Company manufactures a product for which the following standards have been set:

Standard Quantity
or Hours
Standard Price
or Rate
Standard
Cost
Direct materials 3 feet $ 5 per foot $ 15
Direct labor ? hours ? per hour ?

During March, the company purchased direct materials at a cost of $57,090, all of which were used in the production of 3,400 units of product. In addition, 5,200 direct labor-hours were worked on the product during the month. The cost of this labor time was $54,600. The following variances have been computed for the month:

Materials quantity variance $ 900 U
Labor spending variance $ 3,600

U

Labor efficiency variance $ 1,000

U

Required:

1. For direct materials:

a. Compute the actual cost per foot of materials for March.

b. Compute the price variance and the spending variance.

2. For direct labor:

a. Compute the standard direct labor rate per hour.

b. Compute the standard hours allowed for the month’s production.

c. Compute the standard hours allowed per unit of product.
Answer #1
Computation of Standard Direct material cost allowed for actual output :-
a Actual output 3400 units
b Standard Direct materials quantity per unit 3 Feet
c Standard Direct materials quantity for actual output (a*b) 10200 Feet
d Standard Direct materials price per foot $        5.00 per foot
e Standard Direct materials cost for Standard Direct materials Quantity for actual output (c*d) $   51,000
1.a .Computation of actual cost per foot of materials for March :-
a Standard Direct materials cost for Standard Direct materials Quantity for actual output (c*d) $    51,000
b Materials quantity variance $         900 Unfavorable
c Standard Direct materials cost for actual Direct materials quantity purchased and used $    51,900
d Standard Direct materials price per foot $        5.00 per foot
d Actual Direct material quantity purchased and used (c/d)        10,380 Feet
e Actual Direct material Purchase Cost $    57,090
f Actual cost per foot of materials (e/d) $        5.50 per foot
1.b .Computation of price variance and the spending variance :-
a Actual Direct material Purchase Cost $    57,090
b Standard Direct materials cost for actual Direct materials quantity purchased and used $    51,900
c Price variance and the spending variance (a-b) $      5,190 Unfavorable
Verification of variances :-
a.
     Direct materials price variance :-
Direct materials price variance = Actual quantity purchased × ( Actual Price - Standard Price )
=                                                  10,380 × ( 5.50 - 5.00 )
= 10380 × 0.50
= $5,190 Unfavorable    
      Direct materials quantity variance :-
Direct materials quantity variance = Standard Price × ( Actual quantity - Standard quantity )
= 5.00 × ( 10380 - 10200 )
= 5.00 × 180 -
= $900 Unfavorable    
2.a. Computation of the standard direct labor rate per hour :-
a Actual labor cost for Actual hours used $   54,600
b Actual labor hours worked 5200 hours
c Actual direct labor rate per hour (a/b) $      10.50
d Labor spending variance $      3,600 Unfavorable
e Standard labor cost for Actual hours used   (a-d) $   51,000
f Standard labor rate per hour (e/b) $        9.81 per hour
2.b. Computation of the standard hours allowed for the month’s production :-
a Standard labor cost for Actual hours used $   51,000
b Labor efficiency variance $      1,000 Unfavorable
c Standard labor cost for Standard hours for actual output (a-b) $   50,000
d Standard labor rate per hour $        9.81 per hour
e Standard hours allowed for the month’s production (c/d)          5,098 hours s
2.b. Computation of the standard hours allowed per unit of product :-
a Standard hours allowed for the month’s production $      5,098
b Actual output          3,400 units
c Standard hours allowed per unit of product (a/b)            1.50 per hour
Verification of answer
     Direct labor Spending variance :-
Direct labor rate variance = Actual labor hours × ( Actual Rate - Standard Rate )
=                                                    5,200 × ( $      10.50 - $      9.81 )
=                                                    5,200 × 0.69
= $3,600 Unfavorable    
   Direct labor efficiency Variance :-
Direct labor Efficiency Variance = Standard Rate × ( Actual labor hours - Standard labor hours )
= 9.81 × (          5,200 - 5098 )
= 9.81 × 101.96
= $1,000 Unfavorable    

Feel free to ask any clarification, if required. Please provide feedback by thumbs up, if satisfied. It will be highly appreciated. Thank you.